# DCA-Bench: A Benchmark for Dataset Curation Agents

### v1

@@ Abstract
The quality of datasets plays an increasingly crucial role in the research and development of modern artificial intelligence (AI). Despite the proliferation of open dataset platforms nowadays, data quality issues, such as insufficient documentation, inaccurate annotations, and ethical concerns, remain common in datasets widely used in AI. Furthermore, these issues are often subtle and difficult to be detected by rule-based scripts, requiring expensive manual identification and verification by dataset users or maintainers. With the increasing capability of large language models (LLMs), it is promising to streamline the curation of datasets with LLM agents. In this work, as the initial step towards this goal, we propose a dataset curation agent benchmark, DCA-Bench, to measure LLM agents’ capability of detecting hidden dataset quality issues. Specifically, we collect diverse real-world dataset quality issues from eight open dataset platforms as a testbed. Additionally, to establish an automatic pipeline for evaluating the success of LLM agents, which requires a nuanced understanding of the agent outputs, we implement a dedicated Evaluator using another LLM agent. We demonstrate that the LLM-based Evaluator empirically aligns well with human evaluation, allowing reliable automatic evaluation on the proposed benchmark. We further conduct experiments on several baseline LLM agents on the proposed benchmark and demonstrate the complexity of the task, indicating that applying LLMs to real-world dataset curation still requires further in-depth exploration and innovation. Finally, the proposed benchmark can also serve as a testbed for measuring the capability of LLMs in problem discovery rather than just problem-solving. The benchmark suite is available at https://github.com/TRAIS-Lab/dca-bench.
1 Introduction
@@
High-quality datasets have become increasingly crucial for advancing artificial intelligence (AI) [1, 2]. Open datasets platforms, such as Hugging Face [3] or BIG-Bench [4], have substantially accelerated AI research and development by facilitating community contributions. However, community-contributed datasets often encounter subtle data quality issues, including insufficient documentation [5], inaccurate annotations [6], ethical concerns [7], and inconsistency across multiple files (see Appendix A.1 for real-world examples).
@@
There have been existing efforts on standardizing dataset management practices [7, 8, 9], and developing dataset curation toolkits and systems [10, 11]. However, these toolkits and systems often rely heavily on rule-based scripts, which lack the necessary flexibility to detect the aforementioned subtle data quality issues. Consequently, existing techniques are insufficient to tackle the complex challenge of automatically curating datasets at scale on open dataset platforms, where we still primarily rely on dataset users or platform maintainers to identify data quality issues.
@@
Figure 1: Illustration of the instances in DCA-Bench. The Issue Card displays a specific test case in DCA-Bench. Each test case includes essential metadata such as ID, source platform (BIG-Bench in this case), type, tags, hints generated, and other items, as shown in the Issue Card. Relevant dataset files can be found using the Issue ID. Furthermore, DCA-Bench incorporates documentation from the dataset platform along with additional reference materials related to dataset curation and quality. These components are integrated into the API Layer for convenient access. The Curator is asked to detect the issues in files by describing the issue context and pinpointing the location in the file where issues occur. The elements labeled as “title”, “content”, and “involved” are considered as the ground truth, which will be used for evaluating the Curator and is not available as inputs to the Curator.
@@
Recent advancement of large language models (LLMs) has led to promising development of LLM agents for real-world software engineering problems [12]. For certain issues in software development, LLM agents have been shown to be capable of autonomously generating proper code to fix them [13, 14]. Given this development, we envision that LLM agents could become an effective technique for building autonomous dataset curation systems.
@@
However, there is still a significant gap between the need of dataset curation and the state-of-the-art LLM agents for software engineering. Specifically, existing studies primarily focus on developing agents that solve identified and well-defined issues, while dataset curation requires one to discover hidden issues in the community-contributed datasets—a distinct capability different from problem-solving [15, 16] and remaining under-explored for LLM agents.
@@
In response to these challenges, we introduce the Dataset Curation Agents Benchmark (DCA-Bench) as an initial step towards achieving autonomous dataset curation. Specifically, DCA-Bench aims to provide a comprehensive benchmark for evaluating LLM agents’ capability to discover data quality issues across online dataset platforms, the first step of the curation pipeline. Henceforth, we will consistently refer to such an LLM agent as a “Curator” to highlight its role in this task. A well-performed Curator can detect and locate existing issues, which is critical for a follow-up fix by human maintainers or other LLM agents. We collected 91 representative samples from 8 online dataset platforms and classified them into 4 types with 18 tags according to their various content and difficulty. The main features of DCA-Bench include:
@@
  • Real-world Cases with Minimal Simplification: All test cases of DCA-Bench have references to real-world sources, allowing benchmark users to better understand them in the wild. To test the Curator’s ability under the complex real-world environment, DCA-Bench incorporates all relevant dataset files for each test case, not only limited to those flawed ones.
@@
  • Accurate Automatic Evaluation Unlike traditional machine learning tasks, the task of dataset curation does not have labels that can be directly evaluated by scripts. Human-level efforts are required to rate the performance of the Curator, which is not scalable. Therefore, we develop an automatic and accurate evaluation scheme using GPT4 [17] to replace human annotators.
In addition to the benchmark, we implement a straightforward baseline for this task using OpenAI GPT4 Assistant [^1] We use GPT-4-0125-preview with Code Interpreter enabled.. The provided baseline only succeeds in detecting 10.99% issues without hints and 62.64% when given the most specific hint (see Section 3.2), demonstrating the difficulty of this benchmark.
@@
We organize this paper as follows: In Section 2, we introduce the background of the dataset curation problem and existing literature. In Section 3, we go through an overview of the composition of DCA-Bench, the construction process, task definition as well as how we evaluate the performance of the Curator. In Section 4, we carry out experiments to validate the reliability of our automatic evaluation scheme and show preliminary testing results of baseline Curators on DCA-Bench.
2 Related Work
Dataset Quality Management
With the advancement of modern AI, the need for high-quality datasets is surging [1, 2]. Consequently, open dataset platforms like Hugging Face and Kaggle keep growing rapidly. However, recent studies [6, 18, 19] indicate that many popular datasets suffer from errors, biases, or annotation artifacts [20], and lack proper documentation [5]. In addition, common issues in real-world datasets include problems with data loading scripts and file discrepancies, where conflicting information exists between files.
There have been many exisiting efforts on dataset quality management. Gebru et al. [7] outline principles for creating and managing datasets, especially in the context of machine learning. Gong et al. [21] propose several strategies, including dataset profiling, data cleansing, and quality monitoring. More recent works [7, 8, 9] further explore standardizing dataset management practices. Based on these works, researchers have developed several dataset curation toolkits and systems [10, 11]. More recently, Zhou et al. [22] developed a system that is relevant to enhancing data quality. However, they focus on filtering a large-scale LLM-pertaining corpus rather than detecting dataset quality issues. Besides, there are some specialized works focused on detecting harmful content [23, 24] and annotation errors [19, 25]. Nevertheless, those methods are not generalizable enough and lack flexibility to solve challenges such as nuanced ethical biases (see Tab. 8) and incorrect annotations involving logical or factual errors (see Tab. 10). Moreover, studies on file discrepancies(see Tab. 9) and issues remain scarce. Consequently, we still primarily rely on dataset users or platform maintainers to identify these issues in practice.
We envision LLM agents as a promising technique to address the challenges in this space, and propose a benchmark dataset as an initial step towards this goal.
LLMs for Software Engineering
Our work is also relevant to the emerging research direction on using LLMs to solve software engineering tasks. Fan et al. [26] suggest that there are two major categories in this area: (i) code generation and completion, and (ii) maintenance and evolution.
There has been extensive literature on LLMs for code generation and completion, in terms of both models and benchmarks. On the model side, there have been various code generation LLMs [27, 28, 29]. On the benchmark side, datasets have been created to complete functions and short programs from natural language descriptions [12, 30, 31, 32, 33, 34] and retrieve relevant code [35]. We note that the main purpose of the proposed DCA-Bench is to measure LLM agents’ capability of discovering hidden dataset issues instead of generating code, although generating and running some test code may be helpful.
For the second category, maintenance and evolution, it involves tasks such as localizing and fixing bugs [36, 37], improving program performance [38, 39], and refactoring code without changing program behavior [40]. Our work falls into this category while differing from previous studies in two key aspects. Firstly, while most literature focuses on tasks with well-defined outcomes, such as fixing a buggy function, our work addresses non-standard issues like ethical concerns in data points or cross-file inconsistencies. Secondly, few studies have explored using LLMs to maintain dataset repositories, which is increasingly important for the machine learning community.
LLMs as A Proxy of Human Evaluation
Recent studies have shown that LLMs, when carefully prompted, can serve as a good proxy of human evaluations for a number of scenarios, such as evaluating text generation quality [41, 42], reasoning ability [43, 44], and generated image quality [45, 46].
3 The Dataset Curation Agent Benchmark
@@
In this section, we start with an overview of the proposed DCA-Bench in Section 3.1. Following that, we introduce the benchmark test case construction process in Section 3.2 and give a detailed task definition in Section 3.3. In Section 3.4, we introduce our evaluation pipeline in DCA-Bench.
3.1 Benchmark Overview
At a high level, DCA-Bench consists of three main parts: benchmark assets, a benchmark API, and an evaluation pipeline.
Benchmark assets include test cases and reference resources. For test cases, DCA-Bench includes 91 dataset issues from eight dataset platforms. Statistics about the benchmark are shown in Tab. 2. Each test case includes multiple files, providing a minimal environment for identifying hidden issues. We classify those cases according to the number of files and issues involved, and further assign 18 tags to categorize them, which include “data-problem”, “document-problem”, “infrastructure-problem”, “ethical/legal-risk”, and more. Statistics by file and issue count are shown in Tab. 3. The detailed definitions of the types and tags are provided in Appendix A.3. Readers can refer to Appendix A.10 for some concrete examples of dataset issues.
For reference resources, DCA-Bench includes each platform’s documentation and external materials. Depending on each platform, the documentation may consist of the README files, contribution guidelines, and dataset formats, setting out specific rules for contributing to the platform. The external materials encompass research papers on dataset curation, blogs about best practices for dataset maintenance, and various tutorials. These resources offer additional context that might help identify discrepancies between platform requirements and the actual datasets. Benchmark users can utilize this information to improve the performance of their Curators using RAG [47] or other methods.
The benchmark API is a Python library associated with the DCA-Bench, where it loads the benchmark assets and provides formatted inputs to the Curators and the LLM-based Evaluator. Benchmark users can use the API to easily run their Curator on the benchmark.
The key component of the evaluation pipeline is an LLM-Based Evaluator, which is an LLM agent that automatically evaluates the performance of the Curators. The detailed design of the Evaluator is discussed in Section 3.4, and we demonstrate that it empirically aligns well with human evaluations in Section 4.1.
Figure 2: Distribution of source platform in DCA-Bench.
@@ 3.1 Benchmark Overview
Table 1: File statistics of DCA-Bench. #Token is calculated using the tiktoken package to tokenize the file content. Non-text files are skipped. See Section A.4 for a detailed calculating procedure.
#Samples Avg. #File Avg. #Token
91 2.27 2.96 ×$\times 10^{7}$
Table 2: Test case statistics by file and issue count.
Single-File Multi-File
Single-Issue 19 54
Multi-Issue 6 12
Table 3: Statistics of four issue tags in DCA-Bench. Each test case can contain more than one tag. See a full statistics in Section A.3.
data document infrastructure ethical-legal
71 28 13 4
3.2 Test Case Construction
@@
The test cases in DCA-Bench are collected or modified from the real issues reported by users or maintainers of eight dataset platforms, including Hugging Face  [^2] See https://huggingface.co/datasets, BIG-Bench [4], Kaggle, OpenML, TensorFlow Dataset, Open-Data-Registry, Five-Thirty-Eight, and Graph Learning Benchmark (GLI) [48]. The construction of the test cases of the DCA-Bench involved the following two stages.
3.2.1 Issue Collection and Preprocessing
The issue collection and preprocessing stage ensures the selection of relevant and manageable dataset issues. The steps in this stage are summarized as follows:
  1. 1.
@@ 3.2.1 Issue Collection and Preprocessing
    Manual Selection of Dataset Platforms and Issues: We first select dataset platforms where users and maintainers can interact. For example, there is a “Discussion” section for each Kaggle dataset to discuss dataset-related issues, which helps collect meaningful data quality issues. We then manually collect data quality issues from these discussions.
  2. 2.
@@
    Issue Filtering and Modification: We prioritize the issues that have received feedback from dataset maintainers. Otherwise, we manually verify the problems mentioned in the issues to ensure their validity. We exclude issues that are not directly related to dataset files uploaded by contributors (e.g., issues about Tensorflow or Hugging Face data loading APIs). For issues involving multiple sub-issues or file sizes over 512MB, we either discard them or split them into manageable sub-cases with minimal adjustments.
  3. 3.
@@
    File Downloading: After gathering candidate issues, we download the dataset files from the version before the problems are fixed. To simulate real-world scenarios, the dataset files included in each test case are not limited to those with known problems; we also include some additional files in the dataset not directly relevant to the problem.
3.2.2 Hints Generation
@@
Asking Curators to discover hidden issues from raw dataset files can be very challenging. To make it more manageable and test the Curators’ capability in finer granularity, we generate different levels of hints to help the Curators locate the hidden issues. We apply GPT-4 with carefully designed prompts [^3] See Appendix A.6 for the exact prompts. to generate three levels of hints for each test case, in addition to the no-hint setting:
@@ 3.3 Benchmark API
Figure 3: The Task I/O of DCA-Bench. For each test case, the input for the Curator includes dataset files and hints, with reference materials and platform documentation being optional. The Curator is then required to provide a description of the issue and corresponding contextual evidence. The label of the test case includes the issue title, content, the involved file names, and corresponding contextual evidence. Given the output from the Curator and the label, the Evaluator is then asked to rate the performance of the Curator.
3.4 Evaluation Pipeline

### v2

@@ Abstract.
The quality of datasets plays an increasingly crucial role in the research and development of modern artificial intelligence (AI). Despite the proliferation of open dataset platforms nowadays, data quality issues, such as incomplete documentation, inaccurate labels, ethical concerns, and outdated information, remain common in widely used datasets. Furthermore, these issues are often subtle and difficult to be detected by rule-based scripts, therefore requiring identification and verification by dataset users or maintainers–a process that is both time-consuming and prone to human mistakes. With the surging ability of large language models (LLM), it’s promising to streamline the discovery of hidden dataset issues with LLM agents. To achieve this, one significant challenge is enabling LLM agents to detect issues in the wild rather than simply fixing known ones. In this work, we establish a benchmark to measure LLM agent’s ability to tackle this challenge. We carefully curate 221 real-world test cases from eight popular dataset platforms and propose an automatic evaluation framework using GPT-4o. Our proposed framework shows strong empirical alignment with expert evaluations, validated through extensive comparisons with human annotations. Without any hints, most competitive Curator agent can only reveal $\sim$30of the data quality issues in the proposed dataset, highlighting the complexity of this task and indicating that applying LLM agents to real-world dataset curation still requires further in-depth exploration and innovation. The data and code are available at https://github.com/TRAIS-Lab/dca-bench.
Dataset Curation; Large Language Models; Automatic Evaluation
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737422isbn: 979-8-4007-1454-2/2025/08ccs: General and reference General conference proceedingsccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligenceccs: Information systems Data management systems
1. Introduction
@@
High-quality datasets have become increasingly crucial for advancing artificial intelligence (AI) (Jain et al., 2020; Kaplan et al., 2020). Open dataset platforms, such as Hugging Face (Lhoest et al., 2021) and BIG-Bench (Srivastava et al., 2023), have substantially accelerated AI research and development by facilitating community contributions. However, community-contributed datasets often encounter subtle data quality issues, including insufficient documentation (Yang et al., 2024), inaccurate annotations (Klie et al., 2024), and ethical concerns (Gebru et al., 2021).
@@
There have been existing efforts on standardizing dataset management practices (Gebru et al., 2021; Wang et al., 2023; Gan et al., 2024) and developing dataset curation toolkits and systems (Gupta et al., 2021; Mao et al., 2023). However, these toolkits and systems often rely heavily on rule-based scripts, which lack the necessary flexibility to detect the aforementioned subtle data quality issues. Consequently, existing techniques remain insufficient to address the complex challenge of automatically curating datasets at scale on open dataset platforms, where we still primarily rely on dataset users or platform maintainers as dataset curators to identify data quality issues.
@@
Figure 1. Illustration of instances in DCA-Bench. The Issue Card displays a specific test case, including metadata such as Issue ID, source platform (e.g., BIG-Bench), type, associated tags, and hints. Relevant dataset files can be found using the Issue ID. Furthermore, DCA-Bench incorporates documentation from the dataset platform along with additional reference materials related to dataset curation and quality. We provide a convenient API to access data of each test case as well as reference materials. The Curator is asked to detect the issues in files by describing the issue context and pinpointing the location in the file where issues occur. The elements labeled title, content, and involved serve as ground truth for evaluating the Curator’s performance and are hidden from the Curator during testing.
@@
Recent advancement of large language models (LLMs) has led to promising development of LLM agents for real-world software engineering problems (Jimenez et al., 2024a). For certain issues in software development, LLM agents have been shown to be capable of autonomously generating proper code to fix them (Mesh, 2024; Tao et al., 2024). Given this development, we envision that LLM agents could become an effective technique for building autonomous dataset curation systems.
@@
However, there is still a significant gap between the need of dataset curation and the state-of-the-art LLM agents for software engineering. Specifically, existing studies primarily focus on developing agents that solve identified and well-defined issues, while dataset curation requires one to discover hidden issues in the community-contributed datasets—a distinct capability different from problem-solving (Jay, 1996; Wikipedia contributors, 2024) and remaining under-explored.
@@
In response to these challenges, we introduce the Dataset Curation Agents Benchmark (DCA-Bench) as an initial step towards achieving autonomous dataset curation. Fig.1 illustrates the instance structure in DCA-Bench. Specifically, DCA-Bench aims to provide a comprehensive benchmark for evaluating LLM agents’ capability to discover data quality issues across online dataset platforms, the first step of the curation pipeline. Henceforth, we will consistently refer to such an LLM agent as a “Curator” to highlight its role in this task. A well-performed Curator can detect and locate existing issues, which is critical for a follow-up fix by human maintainers or other LLM agents.
Rather than defining the data quality issue beforehand, we adopt a bottom-up methodology for issue collections and classifications. Specifically, we collect 221 diverse and representative data quality issues from 8 online dataset platforms, covering a broad spectrum of problems such as data errors, documentation issues, file discrepancies, and legal/ethical risks. These issues are further categorized into 4 types and detailed with 18 tags, reflecting their varied content and complexity. The main features of DCA-Bench include:
@@
  • Real-world Cases with Minimal Simplification: All test cases of DCA-Bench have references to real-world sources, allowing benchmark users to better understand them in the wild. To test the Curator’s ability under the complex real-world environment, each test case in DCA-Bench not only contains those flawed ones, but also files that without known issues.
@@
  • Accurate Automatic Evaluation: Unlike traditional machine learning tasks, the task of dataset curation does not have labels that can be directly evaluated by scripts. Human-level efforts are required to rate the performance of the Curator, which is not scalable. Therefore, we develop an automatic and accurate evaluation scheme using LLMs to replace human annotators.
In addition to the benchmark, we implement several straightforward baseline Curators for this task, including recent Open AI models (OpenAI, 2024a; OpenAI et al., 2024), Deepseek R1 (DeepSeek-AI et al., 2025a) and V3 (DeepSeek-AI et al., 2025b), and Meta Llama3.3-70B-Instruct (AI, 2024). The most competitive baselines only succeeds in detecting $\sim$ 29.86% issues without hints and 78.28% when given the most specific hint, demonstrating the difficulty of this benchmark.
@@
We organize this paper as follows: In Section 5, we introduce the background of the dataset curation problem and existing literature. In Section 2, we go through an overview of the composition of DCA-Bench, the construction process, task definition, and the automatic evaluation framework. In Section 3, we carry out experiments to validate the reliability of our automatic evaluation scheme and discuss preliminary testing results of baseline Curators on DCA-Bench.
2. The Dataset Curation Agent Benchmark
@@
In this section, we start with an overview of the proposed DCA-Bench in Section 2.1. Next, we introduce the benchmark construction process in Section 2.2 and give a detailed task definition in Section 2.3. Finally, Section 2.4 presents our evaluation framework.
2.1. Dataset Overview
The dataset assets provided in DCA-Bench consist of test cases and reference resources. Tab. 1 shows basic statistics of test cases. Each test case typically contains multiple files, designed to create a minimal environment for uncovering hidden data quality issues. We classify these test cases into four categories based on the number of files included and number of hidden issues in a test case. Additionally, we assign 18 descriptive tags to further label the cases, such as “data-problem”, “document-problem”, “infrastructure-problem”, and “ethical/legal-risk”. Furthermore, DCA-Bench provides reference resources, including related documentation and external materials such as dataset curation tutorials, to help users enhance Curator performance using methods like RAG (Lewis et al., 2021).
@@ 2.1. Dataset Overview
Table 1. Basic statistics of DCA-Bench. #Token is calculated using the tiktoken package to tokenize the file content. Non-text files are skipped. See Appendix E.2 for a detailed calculating procedure. For tag-level statistics, each test case can contain more than one tag. The definitions of types, tags and detailed statistics are provided in Appendix A.2, along with examples in Appendix A.2.
Statistic Number
Sample-Level #Samples $221$
Avg. #Files/Sample $2.13$
Avg. #Tokens/Sample $3.58\times 10^{6}$
Type-Level Single-Issue Single-File $61$
Single-Issue Multi-File $100$
Multi-Issue Single-File $14$
Multi-Issue Multi-File $46$
Tag-Level data-problem $197$
document-problem $83$
infrastructure-problem $19$
ethical/legal-risk $10$
Table 2. The number of tags in DCA-Bench. Note that the sum of the sub-category might not equal the parent category. For instance, if there’s an ethical or legal risk identified in a dataset document, it should be tagged as [“document-problem”, “ethical/legal-risk”], without including any sub-categories of ”document-problem” that are not relevant.
Category Number Sub-category Number
typo 18
wrong-format 14
inappropriate-file 4
ethical/legal-risk 10
internal-discrepancy 21
cross-file-discrepancy 44
data-problem 197 wrong-value 71
missing-value 15
data-leakage 2
apparent-corruption 40
hidden-corruption 59
document-problem 83 wrong-info 27
insufficient-info 52
infrastructure-problem 19 data-access 4
script-code 15
2.2. Test Case Construction
@@
The test cases in DCA-Bench are collected from the real issues reported by users or maintainers of eight dataset platforms, including Hugging Face Dataset (Lhoest et al., 2021), BIG-Bench (Srivastava et al., 2023), Kaggle (Kaggle, [n. d.]), OpenML (Vanschoren et al., 2014), TensorFlow Dataset (Abadi et al., 2015), Open-Data-Registry (AWS-Labs, 2024), Five-Thirty-Eight (FiveThirtyEight, 2024), and Graph Learning Benchmark (GLI) (Ma et al., 2022). The construction of the test cases of the DCA-Bench involved the following two stages.
Issue Collection and Preprocessing
At this stage, we select and preprocess relevant and manageable data quality issues with the following steps:
  1. (1)
@@
    Selection of Dataset Platforms and Issues: We focus on dataset platforms where users and maintainers can interact. For example, there is a “Discussion” section in each Kaggle dataset for discussing dataset-related issues, which helps collect meaningful data quality issues. We then manually collect data quality issues from these discussions. Please refer to A.3 for more details about the choice of platforms and case selections.
  2. (2)
@@
    Issue Filtering and Modification: We prioritize the issues that have received feedback from dataset maintainers. Otherwise, we manually verify the problems mentioned in the issues to ensure their validity. We exclude issues that are not directly related to dataset files uploaded by contributors (e.g., issues about Tensorflow or Hugging Face data loading APIs). For issues involving multiple sub-issues or file sizes over 512MB, we either discard them or split them into manageable sub-cases with adjustments.
  3. (3)
@@
    File Downloading: After gathering candidate issues, we download the dataset files from the version before the problems are fixed. In order to simulate real-world scenarios, the dataset files included in each test case are not limited to those involving issues.
Hints Generation
@@
Asking Curators to discover hidden issues from raw dataset files can be very challenging. To make it more manageable and test the Curators’ capability in finer granularity, we generate different levels of hints to help the Curators locate the hidden issues. We apply GPT-4 with carefully designed prompts [^1] See Appendix B.8 for prompts used for the Hint Writer. to generate three levels of hints for each test case, in addition to the no-hint setting:
@@ 2.3. Benchmark API
Figure 2. The Task I/O of DCA-Bench. For each test case, the input for the Curator includes dataset files and hints, with reference materials and platform documentation being optional. The Curator is then required to provide a description of the issue and corresponding contextual evidence. The label of the test case includes the issue title, content, the involved file names, and corresponding contextual evidence. Given the output from the Curator and the label, the Evaluator is then asked to rate the performance of the Curator.
2.4. Evaluation Pipeline